Use the quadratic formula to solve the equation.
–4x2 – 3x + 2 = 0

Answer by sachi(548)

–4x2 – 3x + 2 = 0
or 4x2 + 3x - 2 = 0 (multiplying - on both sides)
or 16x2 +12x -8 = 0 (multiplying 4 on both sides)
or (4x)2 +2*4x*3/2+(3/2)2-(3/2)2-8=0
or (4x+3/2)2=(3/2)2+8=9/4+8=41/4
or 4x+3=sqrt(41/4)=(sqrt41)/2
or 4x=-3+(sqrt41)/2
or x=-3/4+(sqrt41)/8
ans